# Simulating conditioned diffusions on manifolds

### v2

@@ 6. Other types of conditionings
This corresponds to a conditioning where $X$ started in $X_{0}=x_{0}$ and forced into the marginal distribution with density $q$ at time $T$, such as seen in in Baudoin, (2002). This is important for generative diffusion models on manifolds.
@@ 7. Proofs
    For all $k$,
    $\exp\left(\int_{0}^{t}\frac{{\mathfrak{A}}g(s,X_{s})}{g(s,X_{s})}\,\mathrm{d}s\right)\mathbf{1}_{A_{k}^{t}}$
    is ${\mathbb{P}}^{g}$-almost surely uniformly bounded in $t$ on $[t_{0},T)$.
@@
Here, the first equality follows from $E_{t}^{\Phi_{*}h}(Y)=E_{t}^{h}(X)$, which was derived in the proof of 4.1. The final equality follows since ${\mathbb{Q}}^{\Phi_{*}h}$ is the pushforward measure ${\mathbb{P}}^{h}$ under $\Phi$. To see that Assumption (7.1d) is satisfied, note that from the proof of 4.1, we obtain ${\mathfrak{A}}_{Y}(\Phi_{*}g)(s,Y_{s})=({\mathfrak{A}}_{X}g)(s,X_{s})$ and thus
@@
$\left\lvert\nabla\log\kappa(t,x,y)\right\rvert\leq C\left(\frac{{\mathrm{dist}}(x,x_{T})}{t}+\frac{1}{\sqrt{t}}\right)$
@@
$\lim_{t\uparrow T}{\mathbb{E}}\left(g(t,X_{t})\mid X_{s}=x\right)$ $=\lim_{t\uparrow T}\int\kappa({T-t},y,x_{T})p(s,x;t,y)\,\mathrm{d}\mathrm{Vol}_{M}(y)$ (7.8)
$=p(s,x;T,x_{T})=h(s,x)$
@@
where $\gamma_{6}=(2d-1)/2+d/(2\bar{\gamma})$. The function of $\tau$ over which the $\sup$ is taken is concave and attains it’s maximum at $\left(\gamma_{6}\bar{\gamma}/(k\varepsilon)\right)^{-1/\varepsilon}$. Hence,
@@
Since the above term is integrable on $[0,T]$, (7.1d) is satisfied. We proof (7.1e) separately in 7.6. The result now follows from 7.1, where the form of the Radon-Nikodym derivative in Equation 7.1 follows from the observation that ${\mathfrak{A}}g/g=V(\log g)$.
@@
$=\lim_{t\uparrow T}{\mathbb{E}}\left(\frac{f_{s}(X)}{h(0,x_{0})}{\mathbb{E}}\left(g(t,X_{t})\mid\mathcal{F}_{s}\right)\right)$
$=={\mathbb{E}}\left(f_{s}(X)\frac{h(s,X_{s})}{h(0,x_{0})}\right)={\mathbb{E}}^{h}f_{s}(X)$
@@
where $M^{\xi}$ denotes the process
$M^{\xi}_{t}=\int_{t_{0}}^{t}\eta(s)\nabla\log g_{s}\,\mathrm{d}W_{s}$
@@
Moreover, we infer from (7.5) that
${\mathrm{dist}}(X_{t},x_{T})\leq\sqrt{\frac{2(T-t)\xi_{t}}{\eta(t)}}$
Hence, by 7.4, there is a positive constant $c_{1}$ such that
@@
$M_{t}^{\xi}\leq C\log\left(\frac{1}{T-t}\right)+\frac{c_{1}^{2}}{2}\int_{t_{0}}^{t}\eta(s)^{2}\left(\sqrt{\frac{2\xi_{s}}{\eta(s)(T-s)}}+\frac{1}{\sqrt{T-s}}\right)\,\mathrm{d}s$ (7.17)
@@
In order to compute ${\mathfrak{A}}^{g}$, we apply the identity
This yields
${\mathfrak{A}}^{g}\zeta_{t}$ $=-\log g_{t}\partial_{t}\eta(t)-\eta(t)\partial_{t}\log g_{t}-\eta(t){\mathfrak{L}}^{g}(\log g)_{t}$ (7.19)
$=\zeta_{t}\partial_{t}\log\eta(t)-\eta(t)\left(\partial_{t}+V+\nabla\log g+\frac{1}{2}\Delta\right)(\log g)_{t}$
$=\zeta_{t}\partial_{t}\log\eta(t)-\eta(t)V(\log g)t-\eta(t)\left\lvert\nabla\log g_{t}\right\rvert^{2}-\eta(t)\left(\partial_{t}+\frac{1}{2}\Delta\right)(\log g)_{t}$
$=\zeta_{t}\partial_{t}\log\eta(t)-\eta(t)V(\log g)_{t}-\frac{1}{2}\eta(t)\left\lvert\nabla\log g_{t}\right\rvert^{2}$
@@
${\mathfrak{A}}^{g}\xi_{t}=\xi_{t}\partial_{t}\log\eta(t)+\eta(t)\frac{2d-1}{2(T-t)}-\eta(t)V(\log g)_{t}-\frac{1}{2}\eta(t)\left\lvert\nabla\log g_{t}\right\rvert^{2}$ (7.20)
@@
$\left\lvert V(\log g)_{t}\right\rvert\leq c_{2}\left(\frac{{\mathrm{dist}}(X_{t},x_{T})}{T-t}+\frac{1}{\sqrt{T-t}}\right)\leq c_{2}\left(\sqrt{\frac{2\xi_{t}}{\eta(t)(T-t)}}+\frac{1}{\sqrt{T-t}}\right)$ (7.22)
Where the second inequality follows from (7.16).
Using the preceding, we can find an upper bound to for $\xi_{t}$, as displayed in (7.13), as follows. We bound $M_{t}^{\xi}$ using (7.17). To bound $\int_{t_{0}}^{t}{\mathfrak{A}}^{g}\xi_{s}\,\mathrm{d}s$, we use the expression derived in (7.21), in which we apply (7.22). By combining all of this, we obtain
@@
$\xi_{t}\leq\xi_{t_{0}}+C\log\left(\frac{1}{T-t}\right)+\int_{t_{0}}^{t}\ell_{0}(s)\,\mathrm{d}s+\int_{t_{0}}^{t}\ell_{1}(s)\xi_{2}^{1/2}\,\mathrm{d}s+\int_{t_{0}}^{t}\ell_{2}(s)\xi_{2}\,\mathrm{d}s,$ (7.24)
@@
By observation (ii), the map $s\mapsto\eta(s)/(T-s)$ is integrable on $[t_{0},T)$. Hence, $\int_{t_{0}}^{t}\ell_{0}(s)\,\mathrm{d}s$ and $\int_{t_{0}}^{t}\ell_{1}(s)\,\mathrm{d}s$ are bounded in $\lim_{t\uparrow T}$. Moreover, since
@@
and, by observation (iii) $\eta(t)\log\left(\frac{1}{T-t}\right)$ is bounded, $\eta(t)\log\left(\frac{1}{T-t}\right)$ is bounded in $t$, $\xi_{t}$ is almost surely bounded. ∎

### v3

@@ 6. Other types of conditionings
This corresponds to a conditioning where $X$ is started at $X_{0}=x_{0}$ and forced into the marginal distribution with density $q$ at time $T$, such as seen in Baudoin, (2002). This is important for generative diffusion models on manifolds.
@@ 7. Proofs
    For all $k$, $\Psi_{T,g}(X)\mathbf{1}_{A_{k}^{t}}$, with $\Psi_{T,g}(X)$ defined in (3.9), is ${\mathbb{P}}^{g}$-almost surely uniformly bounded in $t$ on $[t_{0},T)$.
@@
Here, the first equality follows from $E_{t}^{\Phi_{*}h}(Y)=E_{t}^{h}(X)$, which was derived in the proof of 4.1. The final equality follows from Assumption (7.1c).
To see that Assumption (7.1d) is satisfied, note that from the proof of 4.1, we obtain ${\mathfrak{A}}_{Y}(\Phi_{*}g)(s,Y_{s})=({\mathfrak{A}}_{X}g)(s,X_{s})$ and thus
@@
$\left\lvert\nabla\log\kappa(t,x,y)\right\rvert\leq C\left(\frac{{\mathrm{dist}}(x,y)}{t}+\frac{1}{\sqrt{t}}\right)$
@@
$\lim_{t\uparrow T}{\mathbb{E}}\left(g(t,X_{t})\mid X_{s}=x\right)$ $=\lim_{t\uparrow T}\int\kappa({T-t},y,x_{T})p(s,x;t,y)\,\mathrm{d}\mathrm{Vol}_{M}(y)=p(s,x;T,x_{T})=h(s,x)$ (7.8)
@@
where $\gamma_{6}=(2d-1)/2+d/(2\bar{\gamma})$. The function of $\tau$ over which the $\sup$ is taken is concave and attains its maximum at $\left(\gamma_{6}\bar{\gamma}/(k\varepsilon)\right)^{-1/\varepsilon}$. Hence,
@@
Since $\beta_{1}(t)$ is bounded on $[t_{0},T)$ (by (iii)) and tends to 00 as $t\uparrow T$, the above term is integrable on $[t_{0},T]$, (7.1d) is satisfied. We proof (7.1e) separately in 7.6. The result now follows from 7.1, where the form of the Radon-Nikodym derivative in Equation 7.1 follows from the observation that ${\mathfrak{A}}g/g=V(\log g)$.
@@
$=\lim_{t\uparrow T}{\mathbb{E}}\left(\frac{f_{s}(X)}{h(0,x_{0})}{\mathbb{E}}\left(g(t,X_{t})\mid\mathcal{F}_{s}\right)\right)={\mathbb{E}}\left(f_{s}(X)\frac{h(s,X_{s})}{h(0,x_{0})}\right)={\mathbb{E}}^{h}f_{s}(X)$
@@
where $M^{\xi}$ denotes the process $M^{\xi}_{t}=\int_{t_{0}}^{t}\eta(s)\nabla\log g_{s}\,\mathrm{d}W_{s}$.
@@
Moreover, we infer from (7.5) that ${\mathrm{dist}}(X_{t},x_{T})\leq\sqrt{{2(T-t)\xi_{t}}/{\eta(t)}}$. Hence, by 7.4, there is a positive constant $c_{1}$ such that
@@
$M_{t}^{\xi}\leq C\log\left(\frac{1}{T-t}\right)+\frac{c_{1}^{2}}{2}\int_{t_{0}}^{t}\eta(s)^{2}\left(\sqrt{\frac{2\xi_{s}}{\eta(s)(T-s)}}+\frac{1}{\sqrt{T-s}}\right)^{2}\,\mathrm{d}s$ (7.17)
@@
In order to compute ${\mathfrak{A}}^{g}\zeta_{t}$, we use the identity from 7.7. This yields
${\mathfrak{A}}^{g}\zeta_{t}$ $=-\log g_{t}\partial_{t}\eta(t)-\eta(t)\partial_{t}\log g_{t}-\eta(t){\mathfrak{L}}^{g}\log g_{t}$ (7.19)
$=\zeta_{t}\partial_{t}\log\eta(t)-\eta(t)\left(\partial_{t}+V+\nabla\log g+\frac{1}{2}\Delta\right)\log g_{t}$
$=\zeta_{t}\partial_{t}\log\eta(t)-\eta(t)V(\log g)t-\eta(t)\left\lvert\nabla\log g_{t}\right\rvert^{2}-\eta(t)\left(\partial_{t}+\frac{1}{2}\Delta\right)\log g_{t}$
$=\zeta_{t}\partial_{t}\log\eta(t)-\eta(t)V\log g_{t}-\frac{1}{2}\eta(t)\left\lvert\nabla\log g_{t}\right\rvert^{2}$
@@
${\mathfrak{A}}^{g}\xi_{t}=\xi_{t}\partial_{t}\log\eta(t)+\eta(t)\frac{2d-1}{2(T-t)}-\eta(t)V\log g_{t}-\frac{1}{2}\eta(t)\left\lvert\nabla\log g_{t}\right\rvert^{2}$ (7.20)
@@
$\left\lvert V\log g_{t}\right\rvert\leq c_{2}\left(\frac{{\mathrm{dist}}(X_{t},x_{T})}{T-t}+\frac{1}{\sqrt{T-t}}\right)\leq c_{2}\left(\sqrt{\frac{2\xi_{t}}{\eta(t)(T-t)}}+\frac{1}{\sqrt{T-t}}\right)$ (7.22)
Where the second inequality follows again follows from (7.5) as ${\mathrm{dist}}(X_{t},x_{T})\leq\sqrt{{2(T-t)\xi_{t}}/{\eta(t)}}$
Using the preceding, we can find an upper bound for $\xi_{t}$, as displayed in (7.13), as follows. We bound $M_{t}^{\xi}$ using (7.17). To bound $\int_{t_{0}}^{t}{\mathfrak{A}}^{g}\xi_{s}\,\mathrm{d}s$, we use the expression derived in (7.21), in which we apply (7.22). By combining all of this, we obtain
@@
$\xi_{t}\leq\xi_{t_{0}}+C\log\left(\frac{1}{T-t}\right)+\int_{t_{0}}^{t}\ell_{0}(s)\,\mathrm{d}s+\int_{t_{0}}^{t}\ell_{1}(s)\xi_{s}^{1/2}\,\mathrm{d}s+\int_{t_{0}}^{t}\ell_{2}(s)\xi_{s}\,\mathrm{d}s,$ (7.24)
@@
By observation (ii), the map $s\mapsto\eta(s)/(T-s)$ is integrable on $[t_{0},T)$ and thus $\ell_{0}$ and $\ell_{1}$ are integrable as well as linear combinations of compositions that preserve integrability. Hence, $\int_{t_{0}}^{t}\ell_{0}(s)\,\mathrm{d}s$ and $\int_{t_{0}}^{t}\ell_{1}(s)\,\mathrm{d}s$ are bounded in $\lim_{t\uparrow T}$. Moreover, since
@@
and, by observation (iii) $\eta(t)\log\left(\frac{1}{T-t}\right)$ is bounded, $\xi_{t}$ is almost surely bounded. ∎
Lemma 7.7.
Suppose $g$ satisfies the heat equation $(\partial_{t}+\frac{1}{2}\Delta)g=0$ Then
Proof.
Using the product rule for the divergence operator
$\frac{1}{2}\Delta\log g$ $=\frac{1}{2}\mathrm{div}\left(\frac{\nabla g}{g}\right)=\frac{1}{2}\left\langle\nabla\left(\frac{1}{g}\right),\nabla g\right\rangle+\frac{1}{2g}\mathrm{div}(\nabla g)$
$=-\left\langle\frac{\nabla g}{g},\frac{\nabla g}{g}\right\rangle-\frac{\partial_{t}g}{g}=-\left\lvert\nabla\log g\right\rvert^{2}-\partial_{t}\log g$
@@
Appendix B Numerical approximation of the guiding term on the Poincaré disk
To evaluate the guiding term of the diffusion bridges, we note that
$\nabla_{x}\log p(s,x;t,y)=\nabla_{x}\log\mathbb{E}_{N(0,t-s)}\left[\frac{Z}{\sqrt{\cosh Z-\cosh\rho_{x}}}\mathbf{1}{\{Z>\rho_{x}\}}\right].$
Hence, the guiding term can be estimated using Monte Carlo sampling. Since $\rho_{x}\geq 0$ this will be very inefficient. To resolve this, we propose importance sampling. Note that with $\tau=\sqrt{t-s}$
$\nabla_{x}\log\mathbb{E}_{N(0,\tau^{2})}\left[\frac{Z}{\sqrt{\cosh Z-\cosh\rho_{x}}}\mathbf{1}_{\{Z>\rho_{x}\}}\right]=\nabla_{x}\log{\mathbb{E}}_{N(\mu,\tau^{2})}\left[\frac{Z\exp\left(-\frac{\mu Z}{\tau^{2}}+\frac{\mu^{2}}{2\tau^{2}}\right)}{\sqrt{\cosh Z-\cosh\rho_{x}}}\mathbf{1}\{Z>\rho_{x}\}\right]$
$\qquad\qquad=\nabla_{x}\log{\mathbb{E}}_{N(0,1)}\left[\frac{(\mu+\tau Z)\exp\left(-\frac{\mu Z}{\tau}-\frac{\mu^{2}}{2\tau^{2}}\right)}{\sqrt{\cosh(\mu+\tau Z)-\cosh\rho_{x}}}\mathbf{1}\{\mu+\tau Z>\rho_{x}\}\right]$
$\qquad\qquad=-\frac{1}{2\tau^{2}}\nabla_{x}\mu^{2}+\nabla_{x}\log{\mathbb{E}}_{N(0,1)}\left[\frac{(\mu+\tau Z)\exp\left(-\frac{\mu Z}{\tau}\right)}{\sqrt{\cosh(\mu+\tau Z)-\cosh\rho_{x}}}\mathbf{1}\{\mu+\tau Z>\rho_{x}\}\right].$
To ensure that at least $97\%$ of the samples exceed $\rho_{x}$, we choose $\mu\equiv\mu(x,\tau)=\rho_{x}+2\tau$. This implies that the event in the indicator in the last line of the preceding display equals $\{Z>-2\}$. We approximate the term with the expectation by Monte-Carlo simulation. Let $\{Z_{i}\}$ be independent $N(0,1)$-distributed random variables and define
$w(x,\tau,z)=\frac{(\mu(x,\tau)+\tau z)\exp(-\mu(x,\tau)z/\tau)}{\sqrt{\cosh(\mu(x,\tau)+\tau z)-\cosh(\rho_{x})}}\mathbf{1}\{z>-2\}.$
Then we approximate $\nabla_{x}\log p(s,x;t,y)$ by
$-\frac{1}{2\tau^{2}}\nabla_{x}\mu(x,\tau)^{2}+\frac{\nabla_{x}\sum_{i}w(x,\tau,Z_{i})}{\sum_{i}w(x,\tau,Z_{i})},$
which is evaluated by (forward) automatic differentiation. When evaluating this gradient for different values of $x$, we fix the realisations of the random variables $\{Z_{i}\}$, where we can simply discard those realisations that are below $-2$.